14. Given a standard normal distribution, find: (For full credit need to draw the curve and shade the area)

a) P(−1.13 < z < 3.11).

b) P (0.21 < z < 0.81).
Answer #1

P(−1.13 < z < 3.11).

= P(Z <3.11 ) - P(Z <-1.13 )

Using z table   

= 0.9991 - 0.1292

probability= 0.8699

b.

P (0.21 < z < 0.81).

P(Z <0.81 ) - P(Z <0.21 )

Using z table   

=0.791 - 0.5832

=0.2078